OpenStudy (anonymous):

Solve V = π r^2h for h

OpenStudy (anonymous):

for h =?

OpenStudy (anonymous):

yess

OpenStudy (anonymous):

What could you do to isolate the 'h'?

OpenStudy (anonymous):

I don't know

OpenStudy (acacia):

V = πr^2h divide by πr^2 both the sides V/πr^2 = πr^2h/πr^2 ==>( πr^2 will get cancelled on right hand side leaving only h) V/πr^2 = h ==> a. do you get it from there?

OpenStudy (anonymous):

yes I get that much

OpenStudy (acacia):

okay so now assume x to be the 1st odd integer, then (x+2) would be the next immediate odd integer. therefore, the larger of two consecutive odd integers (i.e., x +2) is 7 less than twice the smaller (7 - 2x). x + 2 = 2x - 7 if i'm right is that one of your answers?

OpenStudy (anonymous):

yes thanks much!!!!!!!!!!